# A Prototype Hybrid Mode Cavity for Heterodyne Axion Detection

### v1

@@ Abstract
In the heterodyne approach to axion detection, axion dark matter induces transitions between two modes of a microwave cavity, resulting in a parametrically enhanced signal power. We describe the fabrication and characterization of a prototype normal conducting cavity specifically optimized for heterodyne detection. Corrugations on the cavity walls support linearly polarized hybrid modes which maximize the signal power while strongly suppressing noise. We demonstrate tuning mechanisms which allow one mode’s frequency to be scanned across a 4 MHz range, while suppressing cross-coupling noise by at least 80 dB. A future superconducting cavity with identical geometry to our prototype would have the potential to probe orders of magnitude beyond astrophysical bounds.
@@ 1 Introduction
$m_{a}\simeq 4\times 10^{-6}\,\text{e\kern-0.6458ptV}\,\left(\frac{f}{\mathrm{GHz}}\right)$ (1)
and a symmetry breaking scale of $\sim 10^{12}\,\text{G}\text{e\kern-0.6458ptV}$ [6]. However, significantly lower axion masses of $m_{a}\lesssim 10^{-8}\,\text{e\kern-0.6458ptV}$ are motivated by string theory and grand unification [7, 8, 9, 10, 11, 12, 13, 14]. In this regime, a traditional cavity haloscope would become impractically large.
This problem can be circumvented using the heterodyne approach to axion detection, in which a microwave cavity is excited in a “loaded” mode of angular frequency $\omega_{0}$. This causes the axion’s effective current to oscillate at angular frequency $\omega_{0}\pm m_{a}$, which can resonantly drive a “signal” mode of angular frequency $\omega_{1}$ provided that $m_{a}=|\omega_{0}-\omega_{1}|$. A range of lower axion masses can be probed by tuning the small frequency difference between these modes.
The sensitivity of a heterodyne axion experiment depends on the maximum field that can be excited in the loaded mode, the overlap $C_{\text{sig}}$ of the effective current with the signal mode, and the tunability of their frequency difference. In addition, one must mitigate noise due to power leaking from the loaded mode to the signal mode, which can result from the cross-coupling $\epsilon$, or from mode mixing induced by mechanical vibrations, characterized by mechanical form factors $\eta_{p}$. Several of these considerations are unique to the heterodyne approach, and require a dedicated study.
@@
In this work, we describe the design, fabrication, and characterization of a prototype cavity at SLAC whose geometry is optimized for heterodyne axion detection. The cavity, shown in Fig. 1, is fabricated from six rectangular plates. Due to corrugations on the walls, the cavity carries a pair of nearly degenerate, approximately linearly polarized $\mathrm{HE}_{11}$ hybrid modes. Two corrugated endplates shift the modes relative to each other, leading to an almost ideal signal overlap factor $C_{\text{sig}}$ and suppressed values of $\eta_{p}$. Ports for the driving and readout waveguides are placed inside opposite endplates, where the signal and loaded modes are excluded respectively, automatically leading to a strong suppression of $\epsilon$. One endplate can be deformed to tune the mode frequency difference, and rotated to further decrease $\epsilon$.
Since our proof-of-principle cavity is not superconducting, it has limited sensitivity to axion dark matter. However, our measurements demonstrate that this novel cavity design can simultaneously maximize signal and suppress noise, while maintaining a wide tuning range. In addition, our design can be fabricated out of flat plates, which will simplify the transition to a superconducting niobium cavity in future work.
The paper is organized as follows. Sec. 2 reviews the heterodyne approach to axion detection, and defines the form factors that determine its signal and noise. Sec. 3 reviews the theory of hybrid modes, describes the cavity’s design, and shows numeric calculations of the mode profiles, $\epsilon$, and $\eta_{p}$. Sec. 4 describes the mechanical fabrication of the cavity and its tuning mechanism. Sec. 5 presents results from bench measurements of the finished cavity, which corroborate the results of simulations. In particular, we demonstrate suppression of the cross-coupling down to at least $\epsilon\sim 10^{-4}$, and a tuning range of $4\,\mathrm{MHz}$. In Sec. 6, we review related work and discuss potential next steps. Throughout this work we use natural units, $\hbar=c=1$.
@@ 2 Heterodyne Axion Detection
The heterodyne approach to axion detection was first considered in 2010 by Sikivie [15], who focused on the case of $\sim\mathrm{GHz}$ axions, in the regime $\omega_{0}\ll\omega_{1}\approx m_{a}$. In this case, it was found that heterodyne detection did not provide advantages over the traditional cavity haloscope. A decade later, it was realized that for lighter axions, $m_{a}\ll\omega_{0}\approx\omega_{1}$, the heterodyne approach yields a parametrically enhanced signal power compared to alternatives using a static background field. Several works analyzed the potential sensitivity of this approach [16, 17, 18], exploring cavity designs and noise sources. Here we will briefly review the factors that determine the signal and noise, to establish notation for the following sections.
@@ 2.1 Signal Power
where $\mathbf{J}_{\mathrm{eff}}\simeq g_{a\gamma\gamma}\dot{a}\mathbf{B}_{0}$. We define the signal mode lifetime $t_{1}=Q_{1}/\omega_{1}$, and for simplicity assume it is much shorter than the axion coherence time. Then when the axion drives the signal mode on resonance, $m_{a}\simeq|\omega_{0}-\omega_{1}|$, the signal amplitude scales as
@@
In the steady state, the power driven into the signal mode scales as
@@
where we used the fact that the dark matter density scales as $\rho_{{}_{\mathrm{DM}}}\sim\dot{a}^{2}$, we let $b_{0}^{2}=B_{0}^{2}V$ to match the notation of Ref. [16], and defined the signal form factor
@@
We can read off several distinctive properties of the heterodyne approach from Eq. (4). First, for fixed axion density, the signal power is independent of the axion mass. This provides a strong parametric advantage over an approach using a static magnetic field. In that case, the electric field profile near the magnetic field scales as $\tilde{\mathbf{E}}_{1}\sim m_{a}L\,\tilde{\mathbf{B}}_{1}$ in the magnetoquasistatic limit, leading to $C_{\text{sig}}\sim m_{a}L$ and therefore a signal power suppressed by $P_{\text{sig}}\propto(m_{a}L)^{2}$ whenever $m_{a}L\ll 1$. This suppression holds very generally; it is demonstrated for reentrant and dielectric loaded cavities in Ref. [19] and LC circuits in Ref. [20]. By avoiding this penalty, the heterodyne approach starts out with an enhancement of $\sim 10^{6}$ in signal power at $\mathrm{MHz}$ frequencies.
@@ 2.2 Noise Sources
At sufficiently high axion frequencies, the dominant noise sources in a heterodyne axion experiment are thermal noise, set by the temperature $T\simeq 2\,\mathrm{K}$ of the liquid helium bath, and amplifier noise. However, there are additional noise sources proportional to the power $P_{\text{in}}$ input to the cavity, which become increasingly important at lower axion frequencies. This “leakage” occurs either because the waveguides used to drive and readout the cavity couple partially to the wrong mode (“cross-coupling”), or because vibrations directly transfer energy from the loaded mode to the signal mode (“mode mixing”). We consider these effects in the next two subsections. In both cases, the noise power is suppressed both by a geometric form factor, and by a factor depending on the mode separation $m_{a}=|\omega_{0}-\omega_{1}|$, which falls off at higher $m_{a}$.
@@ 2.2.1 Cross-Coupling
The noise power due to cross-coupling is proportional to $P_{\text{in}}\,\epsilon^{2}\,S_{f}(m_{a})$, where $\epsilon$ is a cross-coupling and $S_{f}(m_{a})$ is a power spectral density describing frequency fluctuations. As discussed in Ref. [16], $S_{f}(m_{a})$ arises from both the frequency spread of the oscillator driving the cavity, and from mechanical vibrations which shift the cavity mode frequencies. Since these factors depend on, e.g. the quality of the high power driving oscillator and vibrations in the cavity’s environment, they cannot be directly characterized in our prototype. Thus, we will focus on $\epsilon$, which depends only on the cavity geometry.
To quantify $\epsilon$, let $Q_{i}$ be the loaded quality factor for the loaded ($i=0$) and signal ($i=1)$ mode. The loss rate in each mode is proportional to $1/Q_{i}$, and can be decomposed into an intrinsic loss, and a loss due to the couplings to the driving and readout waveguides,
@@
Here, $Q_{i}^{\text{int}}$ is the quality factor that mode $i$ would have in the absence of the waveguides. Ideally, the driving waveguide couples only to the loaded mode, and the readout waveguide couples only to the signal mode. Thus, the unwanted cross-couplings are defined by
$\epsilon_{\text{d}}=\frac{1/Q_{1\text{d}}}{1/Q_{0\text{d}}},\qquad\epsilon_{\text{r}}=\frac{1/Q_{0\text{r}}}{1/Q_{1\text{r}}}.$ (7)
[…]
Here, we will infer the cross-couplings using microwave scattering parameters $S_{ij}(\omega)$. (For a pedagogical introduction to $S$-parameters, see Ref. [22].) We critically couple both waveguides, $Q_{0\text{d}}\simeq Q_{0}^{\text{int}}$ and $Q_{1\text{r}}\simeq Q_{1}^{\text{int}}$, so that when the cavity is driven with power $P_{\text{in}}$ through the driving waveguide, on resonance with the loaded mode, no power is reflected, $S_{\text{dd}}(\omega_{0})\approx 0$. For $\epsilon_{\text{r}}\ll 1$, a power $\epsilon_{\text{r}}^{2}\,P_{\text{in}}$ exits through the readout waveguide, so we can identify
$\epsilon_{\text{r}}^{2}\simeq|S_{\text{dr}}(\omega_{0})|^{2},\qquad\epsilon_{\text{d}}^{2}\simeq|S_{\text{rd}}(\omega_{1})|^{2},$ (8)
where the second equality follows from similar reasoning. Note that since the cavity is a reciprocal network, we have $S_{\text{rd}}(\omega)=S_{\text{dr}}(\omega)$. Thus, at each input frequency there is only one physical cross-coupling. In a heterodyne experiment, cross-coupling noise is predominantly due to the signal mode being driven through $\epsilon_{\text{d}}$, and the associated noise power is proportional to $\epsilon_{\text{d}}^{2}$.
@@ 2.2.2 Mechanical Mode Mixing
[…]
where $S$ is the cavity’s inner surface, the approximation $\omega_{0}\simeq\omega_{1}$ was made, and $\bm{\xi}_{p}$ is the dimensionless spatial profile of the mechanical mode, normalized so that the average value of $\xi_{p}^{2}$ within the shell of the cavity is $1$. A mechanical mode with $\omega_{p}\simeq m_{a}$ carrying a fraction $U_{p}$ of the total vibrational energy contributes a mechanical mode mixing noise power proportional to $P_{\text{in}}\,\eta_{p}^{2}\,U_{p}$.
@@ 3 Cavity Design
In this section, we discuss the design of our prototype cavity. Sec. 3.1 discusses how the original conceptual design uses the $\text{HE}_{11}$ hybrid mode to enhance the signal. Sec. 3.2 describes the cavity’s final optimized design, and Sec. 3.3 describes its coupling and frequency tuning. Finally, we numerically compute the cross-coupling $\epsilon$ in Sec. 3.4, and form factors for mechanical mode mixing in Sec. 3.5.

### v2

@@ Abstract
In the heterodyne approach to axion detection, axion dark matter induces transitions between two modes of a microwave cavity, resulting in a parametrically enhanced signal power. We describe the fabrication and characterization of a prototype normal conducting cavity specifically optimized for heterodyne detection. Corrugations on the cavity walls support linearly polarized hybrid modes which maximize the signal power while strongly suppressing noise. We demonstrate tuning mechanisms which allow one mode frequency to be scanned across a 4 MHz range, while suppressing cross-coupling noise by at least 80 dB. A future superconducting cavity with identical geometry to our prototype would have the potential to probe orders of magnitude beyond astrophysical bounds.
@@ 1 Introduction
$m_{a}\simeq 4\times 10^{-6}\,\frac{\text{e\kern-0.6458ptV}}{c^{2}}\,\left(\frac{f}{\mathrm{GHz}}\right)$ (1)
and a symmetry breaking scale of $\sim 10^{12}\,\text{G}\text{e\kern-0.6458ptV}$ [31]. However, significantly lower axion masses of $m_{a}\lesssim 10^{-8}\,\text{e\kern-0.6458ptV}/c^{2}$ are motivated by string theory and grand unification [6, 48, 3, 47, 18, 24, 27, 7]. In this regime, a traditional cavity haloscope would become impractically large.
This problem can be circumvented using the heterodyne approach to axion detection, in which a microwave cavity is excited in a “loaded” mode of angular frequency $\omega_{0}$. This causes the axion’s effective current to oscillate at angular frequency $\omega_{0}\pm\omega_{a}$, where $\omega_{a}=m_{a}c^{2}/\hbar$ is the axion’s angular frequency. The effective current can resonantly drive a “signal” mode of angular frequency $\omega_{1}$ provided that $\omega_{a}=|\omega_{0}-\omega_{1}|$. A range of axion masses can be probed by tuning the frequency difference between these modes. The axion signal power is proportional to the energy stored in the loaded mode, which can be maintained at a high value in a superconducting cavity.
This concept could be used to search for axions with sub-Hz to GHz frequencies. However, in addition to the high signal power, there is also noise power due to “leakage” from the loaded mode to the signal mode, which can result from the cross-couplings $\chi_{\text{d},\text{r}}$ (with $\chi_{\text{d}}$ and $\chi_{\text{r}}$ associated with the driving and readout waveguides, respectively) or from mode mixing induced by mechanical vibrations, characterized by mechanical form factors $\eta_{p}$. These noise sources fall rapidly as the axion frequency increases above $\sim$ kHz. In addition, the tuning range of a typical cavity would be of order $\sim$ MHz. Thus, the heterodyne approach is particularly powerful for $\sim$ kHz to $\sim$ MHz axion frequencies, which are the focus of this work.
To optimize a heterodyne experiment, one must design a cavity that maximizes the energy that can be stored in the loaded mode, and the dimensionless overlap $C_{\text{sig}}$ of the effective current with the signal mode. The mode frequencies should be tunable over a wide range, while maintaining low values for $\chi_{\text{d},\text{r}}$ and $\eta_{p}$ to suppress leakage noise. Many of these considerations are unique to the heterodyne approach, and require a dedicated study.
@@
In this work, we describe the design, fabrication, and characterization of a prototype cavity at SLAC whose geometry is optimized for heterodyne axion detection. The cavity, shown in Fig. 1, is fabricated from six rectangular plates. Due to corrugations on the walls, the cavity carries a pair of nearly degenerate, approximately linearly polarized $\mathrm{HE}_{11}$ hybrid modes. Two corrugated endplates shift the modes relative to each other, leading to an almost ideal signal overlap factor $C_{\text{sig}}$ and suppressed values of $\eta_{p}$. Ports for the driving and readout waveguides are placed inside opposite endplates, where the signal and loaded modes are excluded respectively, automatically leading to a strong suppression of $\chi_{\text{d},\text{r}}$. One endplate can be deformed to tune the mode frequency difference, and rotated to further decrease $\chi_{\text{d},\text{r}}$.
Since our proof-of-principle cavity is not superconducting, it has limited sensitivity to axion dark matter, and we do not plan to use it for a dark matter search. However, our simulations and measurements demonstrate that our novel cavity design can simultaneously maximize signal and suppress noise, while maintaining a wide tuning range. In addition, our design can be fabricated out of flat plates, which will simplify the transition to a superconducting niobium cavity in future work.
The paper is organized as follows. Sec. 2 reviews the heterodyne approach to axion detection, and defines the form factors that determine its signal and noise. (This section uses natural units, $\hbar=c=1$.) Sec. 3 reviews the theory of hybrid modes, describes the cavity’s design, and shows numeric calculations of the mode profiles, $\chi_{\text{d},\text{r}}$, and $\eta_{p}$. Sec. 4 describes the mechanical fabrication of the cavity and its tuning mechanism. Sec. 5 presents results from bench measurements of the finished cavity, which corroborate the results of simulations. In particular, we demonstrate suppression of the cross-coupling down to at least $\chi_{\text{d},\text{r}}\sim 10^{-4}$, and a tuning range of $4\,\mathrm{MHz}$. In Sec. 6, we review related work and discuss potential next steps.
@@ 2 Heterodyne Axion Detection
The heterodyne approach to axion detection was first considered in 2010 by Sikivie [44], who focused on the case of $\sim\mathrm{GHz}$ axions, in the regime $\omega_{0}\ll\omega_{1}\approx m_{a}$. In this case, it was found that heterodyne detection did not provide advantages over the traditional cavity haloscope. A decade later, it was realized that for lighter axions, $m_{a}\ll\omega_{0}\approx\omega_{1}$, the heterodyne approach yields a parametrically enhanced signal power compared to alternatives using a static background field. Several works analyzed the potential sensitivity of this approach [9, 35, 10], exploring cavity designs and noise sources. [^1] Earlier, Refs. [30, 52] (with corrections in Refs. [53, 51]) proposed driving two modes of a cavity and measuring frequency shifts induced by axion dark matter. We will call this the “frequency technique”, in contrast to the “power technique” which is the focus of this work. Here we will briefly review the factors that determine the signal and noise, to establish notation for the following sections.
@@ 2.1 Signal Power
where $\mathbf{J}_{\mathrm{eff}}\simeq g_{a\gamma\gamma}\dot{a}\mathbf{B}_{0}$. We define the signal mode lifetime $t_{1}=Q_{1}/\omega_{1}$, and for simplicity assume it is much shorter than the axion coherence time. Then when the axion drives the signal mode on resonance, $m_{a}\simeq|\omega_{0}-\omega_{1}|$, the signal amplitude is
@@
where $\sim$ indicates equality up to a dimensionless constant. In the steady state, the power driven into the signal mode is
@@
where we used the fact that the dark matter density is $\rho_{{}_{\mathrm{DM}}}\sim\dot{a}^{2}$, we let $b_{0}^{2}=B_{0}^{2}V$ to match the notation of Ref. [9], and defined the signal form factor
@@
We can read off several distinctive properties of the heterodyne approach from Eq. (4). First, for fixed axion density, the signal power is independent of the axion mass. This provides a strong parametric advantage over an approach using a static magnetic field. In that case, the electric field profile near the magnetic field scales as $\tilde{\mathbf{E}}_{1}\sim m_{a}L\,\tilde{\mathbf{B}}_{1}$ in the magnetoquasistatic limit, leading to $C_{\text{sig}}\sim m_{a}L$ and therefore a signal power suppressed by $P_{\text{sig}}\propto(m_{a}L)^{2}$ whenever $m_{a}L\ll 1$. This suppression holds very generally, applying to reentrant cavities [22], dielectric loaded cavities [22, 4], and LC circuits [19]. By avoiding this penalty, the heterodyne approach starts out with an enhancement of $\sim 10^{6}$ in signal power at $\mathrm{MHz}$ frequencies.
@@ 2.2 Noise Sources
At sufficiently high axion frequencies, the dominant noise sources in a heterodyne axion experiment are thermal noise, set by the temperature $T\simeq 2\,\mathrm{K}$ of the liquid helium bath, and amplifier noise. (For concreteness, the thermal occupancy of a GHz frequency mode at this temperature is $T/(2\pi\,\mathrm{GHz})\simeq 40$.)
There are also additional noise sources proportional to the power $P_{\text{in}}$ input to the cavity, which become increasingly important at lower axion frequencies. This “leakage” occurs either because the waveguides used to drive and readout the cavity couple partially to the wrong mode (“cross-coupling”), or because vibrations directly transfer energy from the loaded mode to the signal mode (“mode mixing”). We consider these effects in the next two subsections. In both cases, the noise power is suppressed both by a geometric form factor, and by a factor depending on the mode separation $m_{a}=|\omega_{0}-\omega_{1}|$, which falls off at higher $m_{a}$.
@@ 2.2.1 Cross-Coupling
In general, the noise power due to cross-coupling is proportional to the input power $P_{\text{in}}$, a cross-coupling parameter $\chi^{2}$, and to a spectral density which describes how power is converted across the angular frequency separation $m_{a}$. For example, oscillator phase noise contributes a noise power $\sim P_{\text{in}}(\omega_{0}/Q_{0})\,\chi^{2}\,S_{\varphi}(m_{a})$, where $S_{\varphi}$ is the phase noise spectral density, with units of $1/\mathrm{Hz}$. There are also contributions due to oscillator amplitude noise (though we expect it to be subdominant to phase noise at low $m_{a}$), and from mechanical vibrations which shift the cavity mode frequencies; see Ref. [9] for further discussion. Since these quantities depend on, e.g. the quality of the high power driving oscillator and vibrations in the cavity’s environment, they cannot be directly characterized in our prototype. Thus, we will focus on the cross-coupling, which depends only on the cavity geometry.
To quantify it, let $Q_{i}$ be the loaded quality factor for the loaded ($i=0$) and signal ($i=1)$ mode. The loss rate in each mode is proportional to $1/Q_{i}$, and can be decomposed into an intrinsic loss and a loss due to the couplings to the driving and readout waveguides (quantified by external quality factors $Q_{i\text{d}}$ and $Q_{i\text{r}}$), so that
@@
Here, $Q_{i}^{\text{int}}$ is the intrinsic quality factor, i.e. the quality factor that mode $i$ would have in the absence of the waveguides. Ideally, the driving waveguide couples only to the loaded mode, and the readout waveguide couples only to the signal mode. The strength of this coupling is parametrized by the RF coupling strengths $\beta_{0}=Q_{0}^{\text{int}}/Q_{0\text{d}}$ and $\beta_{1}=Q_{1}^{\text{int}}/Q_{1\text{r}}$. There are also unwanted cross-couplings, which we parametrize by
$\chi_{\text{d}}^{2}=\frac{1/Q_{1\text{d}}}{1/Q_{0\text{d}}},\qquad\chi_{\text{r}}^{2}=\frac{1/Q_{0\text{r}}}{1/Q_{1\text{r}}}.$ (7)
One would generically expect the typical cross-coupling $\chi\sim\chi_{\text{d}}\sim\chi_{\text{r}}$ to be roughly set by the fractional precision with which one can mechanically align a waveguide antenna. Stronger suppressions can be achieved using special cavity geometries. For instance, the MAGO collaboration used a two-sphere cavity coupled to a “magic tee” with an adjustable phase shift to achieve $\chi\sim 10^{-7}$, corresponding to $140\,\text{dB}$ of noise reduction [5]. Since $S_{f}(m_{a})$ decreases steeply as the axion mass increases, Ref. [9] estimated [^2] In that work, the quantities $\chi_{\text{d}}$ and $\chi_{\text{r}}$ were denoted as $\epsilon_{1\text{d}}$ and $\epsilon_{0\text{r}}$, respectively. it becomes subdominant to thermal noise for axion frequencies above $\sim\text{kHz}$ for $\chi\sim 10^{-7}$, and above $\sim\text{MHz}$ for $\chi\sim 10^{-5}$.
Here, we will infer the cross-couplings using microwave scattering parameters $S_{ij}(\omega)$. (For a pedagogical introduction to $S$-parameters, see Ref. [40].) We critically couple both waveguides, $Q_{0\text{d}}\simeq Q_{0}^{\text{int}}$ and $Q_{1\text{r}}\simeq Q_{1}^{\text{int}}$, corresponding to $\beta_{0}=\beta_{1}\simeq 1$, so that when the cavity is driven with power $P_{\text{in}}$ through the driving waveguide, on resonance with the loaded mode, no power is reflected, $S_{\text{dd}}(\omega_{0})\approx 0$. For $\chi_{\text{r}}\ll 1$, a power $\chi_{\text{r}}^{2}\,P_{\text{in}}$ exits through the readout waveguide, so we can identify
$\chi_{\text{r}}^{2}\simeq|S_{\text{dr}}(\omega_{0})|^{2},\qquad\chi_{\text{d}}^{2}\simeq|S_{\text{rd}}(\omega_{1})|^{2},$ (8)
where the second equality follows from similar reasoning. Note that since the cavity is a reciprocal network, we have $S_{\text{rd}}(\omega)=S_{\text{dr}}(\omega)$. Thus, at each input frequency there is only one physical cross-coupling. In a heterodyne experiment, cross-coupling noise is predominantly due to the signal mode being driven through $\chi_{\text{d}}$, and the associated noise power is proportional to $\chi_{\text{d}}^{2}$.
@@ 2.2.2 Mechanical Mode Mixing
$\eta_{p}=V^{1/3}\,\bigg|\int_{S}d\mathbf{S}\cdot\bm{\xi}_{p}\,(\tilde{\mathbf{E}}_{0}^{*}\cdot\tilde{\mathbf{E}}_{1}-\tilde{\mathbf{B}}_{0}^{*}\cdot\tilde{\mathbf{B}}_{1})\bigg|\,$, (9)
where $S$ is the cavity’s inner surface, the approximation $\omega_{0}\simeq\omega_{1}$ was made, and $\bm{\xi}_{p}$ is the dimensionless spatial profile of the mechanical mode, normalized so that the average value of $\xi_{p}^{2}$ within the shell of the cavity is $1$. At the $\sim\mathrm{kHz}$ to $\sim\mathrm{MHz}$ frequencies relevant for heterodyne detection, the energy in mechanical modes is largely determined by the spectrum of external vibrations. Explicitly, a mechanical mode with $\omega_{p}\simeq m_{a}$ carrying a fraction $U_{p}$ of the total vibrational energy contributes a mechanical mode mixing noise power proportional to $P_{\text{in}}\,\eta_{p}^{2}\,U_{p}$.
@@ 3 Cavity Design
In this section, we discuss the design of our prototype cavity. Sec. 3.1 discusses how the original conceptual design uses the $\text{HE}_{11}$ hybrid mode to enhance the signal. Sec. 3.2 describes the cavity’s final optimized design, and Sec. 3.3 describes its coupling and frequency tuning. Finally, we numerically compute the cross-coupling $\chi$ in Sec. 3.4, and form factors for mechanical mode mixing in Sec. 3.5.